OpenStudy (anonymous):

Can someone please help me with this problem? Nancy & Harry are lifting a stone statue and moving it to a new location in their garden. Nancy is pushing the statue with a force of 120 newtons (N) at a 60 degree angle with the horizontal while Harry is pulling the statue with a force of 180 N at a 40 degree angle with the horizontal. What is the magnitude of the combined force they exert on the statue? Round answer to the nearest hundredth.

OpenStudy (saifoo.khan):

Drawing will help.

OpenStudy (anonymous):

I've tried, it confuses me even more because I don't know what I'm doing and if I'm doing it right.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

One way to do it is by x and y components.

OpenStudy (anonymous):

Thank you for giving me an illustration :)

OpenStudy (mathstudent55):

120-N force at 60 deg: x-component: 120*cos(60) y-component: 120*sin(60) 180-N force at 40 deg: x-compnent: 180*cos(40) y-component: 180*sin(40) Resultant's sum of components: x-component: 120*cos(60) + 180*cos(40) = 60 + 137.888 = 197.888 y-component: 120*sin(60) + 180*sin(40) = 103.923 + 115.702 = 219.625

OpenStudy (mathstudent55):

 R = sqrt(197.888^2 + 219.625^2) = R = 288.781 Angle = tan^-1(219.625/197.888) = 47.98 deg Answer: 289 N at 48.0 deg